OpenStudy (anonymous):

For the function f(x) = –6x5 + 160x3, determine the interval(s) for which f(x) decreases. A) (-4,4) B) (-4,0) U (4, infinity) C) (-infinity, -4) U (0,4) D) (-infinity, -4) U (4, infinity)

OpenStudy (anonymous):

@texaschic101 can u plz help :)

OpenStudy (texaschic101):

I am sorry, but I am not good at this type of problem

OpenStudy (texaschic101):

@ganeshie8 ..can you help

OpenStudy (anonymous):

He's offline (: but thanks for your time!

OpenStudy (inowalst):

@thomaster

OpenStudy (zpupster):

i can help a little first take derivative, factor out what you can you should end up with x^2-16 factor this and set your terms to equal zero

OpenStudy (abmon98):

Differentiate the function f(x) = –6x^5 + 160x^3 dy/dx=-30x^4+480x^2 Take both 30 and x^2 as a common factor and set the derivative to zero. 30x^2(-x^2+16)=0 30x^2=0 or -x^2+16=0

OpenStudy (zpupster):

graphing it

OpenStudy (anonymous):

\[f(x)=-6x^5+160 x^3\] \[f \prime(x)=-30x^4+480 x^2\] f(x) decreases if \[f \prime(x)<0\] \[-30 x^2(x^2-16)<0\] now \(x^2>0\) hence \((x^2-16)>0\) \(x^2>16\) \[\left| x \right|>4\] x<-4 and x>4 \[(-\infty,-4)\cup (4,\infty)\]

OpenStudy (zpupster):

then state your intervals (inf,-4) (-4,0) (0,4) (4,inf) then pick a number from each of those intervals plug back into f'x (your derivative eqaution and see if it is positive or negative in that interval positive gives increasing negative decreasing then you should see your answer

OpenStudy (triciaal):

additional when you take the 2nd derivative of the function and substitute the critical x-values if f" is negative the function is increasing when f" is positive the function is decreasing

OpenStudy (anonymous):

thank u all soo much idk who to choose as best answer u were all amazing!!!!!!!!!!!! i think the answer is: (as surjithayer said) (-infinity, -4) U (4,infinity)